# Stability Analysis of Thermohaline Convection With a Time-Varying Shear Flow Using the Lyapunov Method

### v1

@@ III-B Numerical simulation
Here, $\text{blkdiag}(\cdot)$ means the block diagonal. This energy $e(t)$ can be connected to the vector norm $\|\boldsymbol{\psi}\|_{2}$ based on the Rayleigh-Ritz theorem:
@@
To obtain the growth rate $\underline{\lambda}$ from numerical simulations, we then perform a least squares fit of $\ln(e)/2=\underline{\lambda}t+B$ for the second half of the time series. We only fit the second half of the time series to avoid fitting the initial transient response. For each wavenumber pair, we conduct 50 simulations with initial conditions for each state variable $\hat{T},\hat{S},\hat{u},\hat{v},\hat{w}$] as a complex normally distributed random variable generated by randn command in MATLAB. We then compute the maximal growth rate from these 50 simulations. Fig. 2 shows sample results of $e(t)$ and the least squares fit associated with wavenumbers $(k,m_{0})=(0.179,3.01\times 10^{-3})$.
@@ III-C Floquet Theory
For a linear time-varying system, we have the solution as
@@
with $\boldsymbol{I}$ as the identity matrix. We can numerically approximate $\boldsymbol{\Phi}(t,t_{0})$ at $t=t_{0}+n\Delta t$ as
$\boldsymbol{\Phi}(t_{0}+n\Delta t,t_{0})=\prod_{j=0}^{n-1}e^{\boldsymbol{A}(j\Delta t+t_{0})\Delta t}.$ (19)
Then, we compute Floquet multiplier $\gamma:=\text{eig}[\boldsymbol{\Phi}(T,0)]$, where $\text{max}\,|\gamma|$ represents the growth of solution norm over one period, i.e., $\frac{\|\boldsymbol{\psi}(T)\|}{\|\boldsymbol{\psi}(0)\|}$. Thus, we have the temporal growth rate predicted from the Floquet theory as $\lambda_{F}=\ln(\text{max}\,|\gamma|)/T$.
@@ IV-1 Lyapunov Method Accurately Predicts the Growth Rate
The growth rate of the linear time-varying system in (5) was computed over initial vertical wavenumbers $m_{0}\in[0,1.5]$ and $x$-direction wavenumbers $k\in[-0.5,0.5]$ with uniformly sampled points $n_{\text{sample}}=126$ in each wavenumber domain. Figs. 3a and 3b show $\overline{\lambda}$ obtained from the Lyapunov method in §III-A with temporal discretization points $n=400$ and $n=800$, respectively. For comparison, Fig. 3c shows $\underline{\lambda}$ obtained from numerical simulations in §III-B, and Fig. 3d shows $\lambda_{F}$ obtained by the Floquet method in §III-C. In Fig. 3, the growth rate results with $\log_{10}(\lambda)<-4$ were truncated to $\log_{10}(\lambda)=-4$, which are shown as dark blue.
@@
The general structure of the unstable parameter regime (“bulbs”) in each panel of Fig. 3 is consistent across these three methods and also consistent with simulation results in Ref. [4]. There is a symmetry across the $k=0$ axis with one main bulb followed by several thinner, elongated, and disconnected parameter regimes of instability. Small initial vertical wavenumbers $m_{0}$ generally resulted in the greatest instability. The width and height of the unstable parameter regime in Fig. 3a ($n=400$) are slightly lower than those shown in Fig. 3b ($n=800$). This indicates that the Lyapunov method requires large enough temporal discretization points $n$ to predict the growth rate accurately. With increasing $n$, the Lyapunov method approached the results of numerical simulations (Fig. 3c) and the Floquet theory (Fig. 3d). We have also observed that the Lyapunov method with drastic under-sampling on the order of $n\sim 10^{1}$ can lose entirely the instability pattern of $\overline{\lambda}(k,m_{0})$.
@@
IV-2 Eigendecomposition of $\boldsymbol{P}(t)$ Provides Insights into Time-dependent Instabilities
Figure 5: Eigendecomposition of $\boldsymbol{P}(t)$ obtained from the Lyapunov method for $(k,m_{0})=(0.179,3\times 10^{-3})$ with $n=2000$. (a) The two largest eigenvalues $\mu_{1}[\boldsymbol{P}(t)]$ and $\mu_{2}[\boldsymbol{P}(t)]$ over time with the background shear flow $A_{U}(t)$. (b) Eigenvector components corresponding to the dominant eigenvalue over time. Note that $u$ and $v$ components are nearly zero.
Here, we perform an eigendecomposition of $\boldsymbol{P}(t)$ obtained from the Lyapunov method to provide insights into the instantaneous growth rate and principal axes of instability. The eigendecomposition $\boldsymbol{P}(t)=\boldsymbol{Q}(t)\boldsymbol{\Lambda}(t)\boldsymbol{Q}(t)^{-1}$ at each time $t$ was performed using the eigenshuffle routine [19] to ensure continuity and consistency of eigenvectors and eigenvalues over time. This eigendecomposition was performed on the $\boldsymbol{P}(t)$ corresponding to $n=2000$ and the wavenumber pair $(k,m_{0})=(0.179,3\times 10^{-3})$. Fig. 5a shows the largest two eigenvalues $\mu_{1}$ and $\mu_{2}$ of $\boldsymbol{P}(t)$, while other eigenvalues (not shown) have much smaller magnitudes. The peak $\mu_{1}[\boldsymbol{P}(t)]$ aligns closely with the peak strength of the background shear flow $A_{U}(t)$, resulting in the time-dependent instability having roughly half the period of the background shear flow. The maximal magnitude of background shear flow occurs at $t=6.283,12.566$, while the peak of $\mu_{1}[\boldsymbol{P}(t)]$ occurs at $t_{\text{peak, 1}}=5.569$ and $t_{\text{peak, 2}}=11.862$. Fig. 5b then shows each component of the eigenvector associated with $\mu_{1}[\boldsymbol{P}(t)]$. The $v$ velocity component is zero for all time since a 2D model was studied (i.e., $l=A_{V}(t)=0$), and the $u$ velocity component was also negligible. As the peak of $\mu_{1}[\boldsymbol{P}(t)]$ is approached, temperature increasingly drives instability, while salinity diminishes marginally. At the steep drop-off of $\mu_{1}[\boldsymbol{P}(t)]$, temperature is the dominant factor, while the salinity will change sign. During non-peak times, salinity almost exclusively influences the instability.
@@ IV-3 Computational Resources Requirement of the Lyapunov Method
We also test the memory usage and running time for the Lyapunov method as summarized in Table III. Here, we parallelize the computation in the $k$ parameter regime and use one computing node with 126 cores and 500 GB of RAM at UConn Storrs High Performance Computing (HPC) facility to generate results in Fig. 3. We take $n_{\text{sample}}=126$ samples in both $k$ and $m_{0}$ parameter regimes. As the number of temporal discretization points $n$ increases, the computational time and memory requirements of the Lyapunov method increase as a result of the increased number of constraints. The CPU hours increase much faster than the memory usage as we are increasing the number of temporal discretization points $n$. Note that the time per $(k,m_{0})$ and the memory per CPU are estimates of the average usage across all CPUs, whether or not they were idle. Because we do not solve the growth rate $\overline{\lambda}$ for the stable case if the constraints in (13) are feasible with $\overline{\lambda}=0$ (i.e., the system is stable), idle CPUs could have occurred if their allotted work contains more stable parameter regimes than other CPUs. We also report the computational resources required for numerical simulations and the Floquet theory in Tab. III. Here, we found numerical simulations (with 50 random initial conditions) require computational time similar to the Lyapunov method with $n\in[200,400]$, while the Floquet theory will be much faster than the Lyapunov method.
@@
TABLE III: HPC performance while evaluating the growth rate for each method corresponding to Fig. 3. The maximal memory usage and CPU hours are for the computation of $n_{\text{sample}}=126$ horizontal wavenumber $k$ and $n_{\text{sample}}=126$ initial vertical wavenumber $m_{0}$. We parallelize the code in the $k$ parameter regime with 126 cores. Performance of numerical simulation is associated with 50 random initial conditions.
Although Floquet theory is much more computationally efficient, it is limited to analyzing linear time-periodic systems. Numerical simulations, in principle, cannot prove stability, but only provide results for certain initial conditions to support the stability prediction. For nonlinear systems, extensive simulations (e.g., 100,000 simulations with random initial conditions [11]) are required to fully characterize the region of attraction. The Lyapunov method, instead, can rigorously certify the region of attraction for nonlinear systems without extensive simulations (e.g., Ref. [15]) and can be generalized to non-periodically time-varying systems (e.g., Ref. [16]). This work serves as a first step to demonstrate that the Lyapunov approach can predict the growth rate consistent with numerical simulations and Floquet theory for linear time-periodic systems. Our ongoing work will generalize this Lyapunov-type framework to analyze nonlinear non-periodically time-varying systems.
@@ V Conclusion and Future Work
We demonstrated that the Lyapunov method is a viable tool to identify instabilities for a linear time-periodic system describing thermohaline convection with a periodically time-varying background shear flow. In the linear matrix inequalities formulation, we include the time derivative of the weighting matrix $\dot{\boldsymbol{P}}(t)$ of the Lyapunov function candidate, which is approximated by the forward Euler method. With enough temporal discretization points, we find that the growth rate predicted by the Lyapunov method or the Floquet theory converges to that obtained by numerical simulations. The eigenvectors of the weighting matrix $\boldsymbol{P}(t)$ are employed to uncover the instantaneous principal direction of instabilities.
We also benchmark the memory and computational time required by the Lyapunov method and how it varies over the number of temporal discretization points, which suggests future directions for improving the computational efficiency of the Lyapunov method. Another future direction is to explore high-order finite difference methods to approximate $\dot{\boldsymbol{P}}(t)$, which might help improve the convergence of the Lyapunov method. The Lyapunov method can be further extended to study the nonlinear stability and systems with non-periodic time dependence; see e.g., Ref. [16].

### v2

@@ III-B Numerical simulation
Here, $\text{blkdiag}(\cdot)$ means the block diagonal. This energy $e(t)$ can be connected to the vector norm $\|\boldsymbol{\psi}\|_{2}$ based on the Rayleigh-Ritz theorem [8]:
@@
To obtain the growth rate $\underline{\lambda}$ from numerical simulations, we then perform a least squares fit of $\ln(e)/2=\underline{\lambda}t+B$ for the second half of the time series. We only fit the second half of the time series to avoid fitting the initial transient response. For unstable regimes, which we focus on, the unstable growing modes will dominate the system response after the initial transient, while transient growth will be important for stable regimes [21]. For each wavenumber pair, we conduct 50 simulations with initial conditions for each state variable $\hat{T},\hat{S},\hat{u},\hat{v},\hat{w}$] as a complex normally distributed random variable generated by randn command in MATLAB. We then compute the maximal growth rate from these 50 simulations. Fig. 2 shows sample results of $e(t)$ and the least squares fit associated with wavenumbers $(k,m_{0})=(0.179,3.01\times 10^{-3})$, where this wavenumber pair is associated with the maximal growth rate [17].
@@ III-C Floquet Theory
For a linear time-varying system in Eq. (5), we have the solution as
@@
with $\boldsymbol{I}$ as the identity matrix. We can numerically approximate $\boldsymbol{\Phi}(t,t_{0})$ at $t=n\Delta t+t_{0}$ as
$\boldsymbol{\Phi}(n\Delta t+t_{0},t_{0})=\prod_{j=0}^{n-1}e^{\boldsymbol{A}(j\Delta t+t_{0})\Delta t}$ (17)
$=$ $e^{\boldsymbol{A}((n-1\Delta t)+t_{0})\Delta t}\cdots e^{\boldsymbol{A}(\Delta t+t_{0})\Delta t}e^{\boldsymbol{A}(t_{0})\Delta t}$ (18)
Then, we compute Floquet multiplier $\gamma:=\text{eig}[\boldsymbol{\Phi}(T+t_{0},t_{0})]$, where $\text{max}\,|\gamma|$ represents the growth of solution norm over one period, i.e., $\frac{\|\boldsymbol{\psi}(T+t_{0})\|}{\|\boldsymbol{\psi}(t_{0})\|}$. Thus, we have the temporal growth rate predicted from the Floquet theory as $\lambda_{F}=\ln(\text{max}\,|\gamma|)/T$.
@@ IV-1 Lyapunov Method Accurately Predicts the Growth Rate
The growth rate of the linear time-varying system in (5) was computed over initial vertical wavenumbers $m_{0}\in[0,1.5]$ and $x$-direction wavenumbers $k\in[-0.5,0.5]$ with uniformly sampled points $n_{\text{sample}}=126$ for each wavenumber domain. Figs. 3a and 3b show $\overline{\lambda}$ obtained from the Lyapunov method in §III-A with temporal discretization points $n=400$ and $n=800$, respectively. For comparison, Fig. 3c shows $\underline{\lambda}$ obtained from numerical simulations in §III-B, and Fig. 3d shows $\lambda_{F}$ obtained by the Floquet method in §III-C. In Fig. 3, the growth rate results with $\log_{10}(\lambda)<-4$ were truncated to $\log_{10}(\lambda)=-4$, which are shown as dark blue.
@@
The general structure of the unstable parameter regime (“bulbs”) in each panel of Fig. 3 is consistent across these three methods and also consistent with simulation results in Ref. [17]. There is a symmetry across the $k=0$ axis with one main bulb followed by several thinner, elongated, and disconnected parameter regimes of instability. Small initial vertical wavenumbers $m_{0}$ generally resulted in the greatest instability. The width and height of the unstable parameter regime in Fig. 3a ($n=400$) are slightly lower than those shown in Fig. 3b ($n=800$). This indicates that the Lyapunov method requires large enough temporal discretization points $n$ to predict the growth rate accurately. With increasing $n$, the Lyapunov method approached the results of numerical simulations (Fig. 3c) and the Floquet theory (Fig. 3d). We have also observed that the Lyapunov method with drastic under-sampling on the order of $n\sim\mathcal{O}(10^{1})$ can lose entirely the instability pattern of $\overline{\lambda}(k,m_{0})$.
@@
IV-2 Eigendecomposition of $\boldsymbol{P}(t)$
Figure 5: Eigendecomposition of $\boldsymbol{P}(t)$ obtained from the Lyapunov method for $(k,m_{0})=(0.179,3\times 10^{-3})$ with $n=2000$. (a) The two largest eigenvalues $\mu_{1}[\boldsymbol{P}(t)]$ and $\mu_{2}[\boldsymbol{P}(t)]$ over time with the base flow $A_{U}(t)$. (b) Eigenvector components corresponding to the $\mu_{1}[\boldsymbol{P}(t)]$ over time. Note that $u$ and $v$ components are nearly zero.
Here, we perform an eigendecomposition of $\boldsymbol{P}(t)$ obtained from the Lyapunov method to uncover the most dangerous disturbance. We compute eigenvalues and eigenvectors of $\boldsymbol{P}(t)$ at each time $\boldsymbol{P}(t)\boldsymbol{q}_{i}(t)=\mu_{i}[\boldsymbol{P}(t)]\,\boldsymbol{q}_{i}(t)$, $i=1,2,\cdots 5$, leading to the eigendecomposition $\boldsymbol{P}(t)=\boldsymbol{Q}(t)\boldsymbol{\Lambda}(t)\boldsymbol{Q}(t)^{*}$ with $\boldsymbol{\Lambda}(t)=\text{diag}(\mu_{1},\mu_{2},\mu_{3},\mu_{4},\mu_{5})$ and a unitary matrix $\boldsymbol{Q}(t)=[\boldsymbol{q}_{1},\boldsymbol{q}_{2},\boldsymbol{q}_{3},\boldsymbol{q}_{4},\boldsymbol{q}_{5}]$. Eigendecomposition of $\boldsymbol{P}(t)$ at each time $t$ was performed using the eigenshuffle routine [4] to ensure continuity and consistency of eigenvectors and eigenvalues over time. This eigendecomposition allows us to write ${V}(\boldsymbol{\psi},t)=\boldsymbol{\psi}^{*}\boldsymbol{P}(t)\boldsymbol{\psi}=\boldsymbol{\psi}^{*}\boldsymbol{Q}(t)\boldsymbol{\Lambda}(t)\boldsymbol{Q}(t)^{*}\boldsymbol{\psi}=\sum_{i=1}^{5}\mu_{i}[\boldsymbol{P}(t)]|\alpha_{i}|^{2}$, where $\boldsymbol{\psi}=\sum_{i=1}^{5}\alpha_{i}\boldsymbol{q}_{i}=\boldsymbol{Q}\boldsymbol{\alpha}$ with $\boldsymbol{\alpha}:=[\alpha_{1},\cdots,\alpha_{5}]^{\text{T}}$. Based on the proof of Theorem 1, we have $V(\boldsymbol{\psi}(t),t)\leq e^{2\overline{\lambda}(t-t_{0})}V(\boldsymbol{\psi}(t_{0}),t_{0})$, and thus, it is most dangerous to reach the upper bound of $V(\boldsymbol{\psi}(t),t)$ when disturbing the system at the moment $t_{m}$ corresponding to the peak of the largest eigenvalue $\mu_{1}[\boldsymbol{P}(t)]$ using the associated eigenvector $\boldsymbol{q}_{1}(t_{m})$. As a result, we interpret eigenvector $\boldsymbol{q}_{1}(t_{m})$ as the most dangerous disturbance.
Fig. 5a shows the largest two eigenvalues $\mu_{1}$ and $\mu_{2}$ of $\boldsymbol{P}(t)$ corresponding to $n=2000$ and the wavenumber pair $(k,m_{0})=(0.179,3\times 10^{-3})$. The peak $\mu_{1}[\boldsymbol{P}(t)]$ aligns closely with the peak strength of the background shear flow $A_{U}(t)$, suggesting that disturbing the flow at this moment will be most dangerous to reach the upper bound of $V$. Fig. 5b then shows each component of the eigenvector $\boldsymbol{q}_{1}(t)$ associated with $\mu_{1}[\boldsymbol{P}(t)]$. The $v$ velocity component is zero for all time since a 2D model was studied (i.e., $l=A_{V}(t)=0$), and the $u$ velocity component was also negligible. At the moment passing through the peak of $\mu_{1}[\boldsymbol{P}(t)]$ ($t\in[4,6.5]$), the magnitude of the temperature mode $T$ increases, while the salinity mode $S$ decreases slowly and then sharply. This suggests that the most dangerous disturbance is dominated by the temperature mode, which is consistent with the physics that the background temperature $\overline{T}=-z$ is destabilizing.
@@ IV-3 Computational Resources Requirement of the Lyapunov Method
We also test the memory usage and running time for the Lyapunov method as summarized in Table III. Here, we parallelize the computation in the $k$ parameter regime and use one computing node with 126 cores and 500 GB of RAM at UConn Storrs High Performance Computing (HPC) facility to generate results in Fig. 3. We take $n_{\text{sample}}=126$ samples in both $k$ and $m_{0}$ parameter regimes. As the number of temporal discretization points $n$ increases, the computational time and memory requirements of the Lyapunov method increase as a result of the increased number of constraints. The CPU hours increase much faster than the memory usage over the number of temporal discretization points $n$. Note that the time per $(k,m_{0})$ and the memory per CPU are estimates of the average usage across all CPUs, whether or not they were idle. Because we do not solve the growth rate $\overline{\lambda}$ for the stable case if the constraints in (11) are feasible with $\overline{\lambda}=0$ (i.e., the system is stable), idle CPUs could have occurred if their allotted work contains more stable parameter regimes than other CPUs. We also report the computational resources required for numerical simulations and the Floquet theory in Tab. III. Numerical simulations (with 50 random initial conditions) require computational time similar to the Lyapunov method with $n\in[200,400]$, while the Floquet theory will be much faster than the Lyapunov method.
@@
TABLE III: HPC performance for each method corresponding to Fig. 3. The maximal memory usage and CPU hours are for the computation of $n_{\text{sample}}=126$ horizontal wavenumber $k$ and $n_{\text{sample}}=126$ initial vertical wavenumber $m_{0}$. We parallelize the code in the $k$ parameter regime with 126 cores. Performance of numerical simulation is associated with 50 random initial conditions.
Although Floquet theory is much more computationally efficient, it is limited to analyzing linear time-periodic systems. The Lyapunov method, instead, can be used to analyze the stability of nonlinear systems [13] and non-periodically time-varying systems [21]. This work serves as a first step to demonstrate that the Lyapunov approach can predict the growth rate consistent with numerical simulations and Floquet theory for linear time-periodic systems.
@@ V Conclusion and Future Work
We demonstrated that the Lyapunov method is a viable tool to identify instabilities for a linear time-periodic fluid system describing thermohaline convection with a periodic time-varying background shear flow. In the linear matrix inequalities formulation, we include the time derivative of the weighting matrix $\dot{\boldsymbol{P}}(t)$ of the Lyapunov function candidate, which is approximated by the forward Euler method. With enough temporal discretization points, we find that the growth rate predicted by the Lyapunov method or the Floquet theory converges to that obtained by numerical simulations. The eigenvectors of the weighting matrix $\boldsymbol{P}(t)$ are employed to uncover the most dangerous disturbances.
We also benchmark the memory and computational time required by the Lyapunov method and how it varies over the number of temporal discretization points, which suggests future directions for improving the computational efficiency of the Lyapunov method. Another future direction is to investigate other high-order numerical methods to approximate $\dot{\boldsymbol{P}}(t)$ in LMI. Moreover, we also aim to leverage this work as a building block to extend the Lyapunov method to analyze the nonlinear stability [13], nonlinear input-output analysis [20], and non-periodically time-varying systems [21].